OpenStudy (anonymous):

find d^2y/dx^2 if xlnx-y=0

hartnn (hartnn):

y=xln x use product rule to first find dy/dx know the product rule ?

OpenStudy (anonymous):

yes i know the product rule, but i though that because the equation was equal 0 i should use implicit differentiation

hartnn (hartnn):

you can use implicit differentiation, in x ln x-y=0 you'll get the same thing.

hartnn (hartnn):

(.....) -dy/dx = 0 (.....)=dy/dx<----implicit xln x =y (....) =dy/dx

OpenStudy (anonymous):

oh I see, thank you very much for your help

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

i see you are new here, so i would : \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

hartnn (hartnn):

if you have any doubts regarding browsing this site, you can aske me :)